OpenStudy (anonymous):

Write a function rule for the cost of movie tickets shown in the table below. Please show your steps to your solution. # of Tickets 2 3 5 Cost in $ 17.50 26.25 43.75

OpenStudy (anonymous):

making a graph of the points will help us see that the information given is a straight line. the final answer we will want will be in the form of y=mx+b

OpenStudy (anonymous):

y=17.5+b

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

close! we have y=mx+b we need to find m, its the slope between two points, do you remember how to do this?

OpenStudy (anonymous):

y2-y1

OpenStudy (anonymous):

y=8.75x+1

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

(yx - yx)/(x2 - x1)

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

now we need to solve for 'b'

OpenStudy (anonymous):

take any point (x, y) plug it into y=8.75x + b and solve for 'b'

OpenStudy (anonymous):

?